# Modelling Lateral Spread in Wire Flat Rolling

### v2

@@ 1 Introduction
Applications including sawblades, springs, piston rings, and transformers depend critically on flattened wires (utsunomiya2001three). These wires are usually made by a flat rolling technique, where a wire with a circular cross-section is cold rolled between cylindrical rolls — sometimes through several passes — to achieve a particular width and thickness (as illustrated in figure 1).
Figure 1: A diagram of wire flat rolling process; the wire initially has a circular cross-section with diameter $d_{0}$ and is flattened to a barrel shape cross section with the lateral spread $W_{t}$ and the contact width $W_{c}$. Adapted from Figure 1 of carlsson1998contact.
@@
In cold rolling of a sheet, deformation in the lateral direction often remains within the elastic range due to the product’s geometry. However, this is not the case when an initial width-to-thickness ratio is less than 6, such as in wide strip or plate, or as low as 1 in round wire (chitkara1966some). In such cases, the plastic flow in the roll gap at the start of rolling schedule is inherently three-dimensional, complicating the analysis. For round wire, this complexity is further increased by the transformation of a round cross-section into a rectangular shape with bulged edges during the first pass.
Although research exists on the modelling of lateral spread in thick plate rolling (e.g. chitkara1966some; lahoti1974hill; oh1975approximate; Nikhaily; Kennedy), wire flat rolling has received less attention. kazeminezhad2008error (kazeminezhad2005width; kazeminezhad2006pressure; kazeminezhad2008error) published a series of papers studying different parameters in wire rolling. kazeminezhad2005width developed a relationship for the width of the contact area between the rolls and the wire. Their equation was based on experimental and FE observations of the formation of X-shaped shear bands in the wire cross-section during the rolling process (Semiatin1984FormabilityAW; pesin2002mathematical), along with the assumption that as the height reduction increases, the shear bands rotate while maintaining a constant length. kazeminezhad2006pressure then applied this equation alongside the slab method, leading to a pressure hill distribution. The study employs the Tresca yield criterion, assuming the longitudinal stress $\sigma_{xx}$ to be the minimum stress, the vertical stress $\sigma_{yy}$ to be the maximum stress, and the compressive stress in the lateral direction, $\sigma_{zz}$, to lie between them. Although $\sigma_{zz}$ is implicitly included in the yield function, and therefore in calculating the roll pressure, the value of $\sigma_{zz}$ is unknown and must be approximated to find the lateral spread. kazeminezhad2005width therefore assumed that $\sigma_{zz}$ takes a value between plane strain and plane stress, written as $\alpha\sigma_{yy}$, where $\alpha$ is a fitting parameter taking values between 0 and 0.5. Using flow rule equations and setting $\sigma_{xx}$ to be zero at the end of the roll gap, they then wrote
@@
where $W_{t}$ is the lateral spread, $d_{0}$ is the initial diameter, and $2h_{1}$ is the final height of the wire. The first use of this formulation has been attributed to hill1955 for plate rolling and has also been derived by several other authors (e.g. chitkara1966some; applied1961formula), who experimentally observed a linear relationship between $\ln(W_{t}/W_{0})$ and $\ln(h_{0}/h_{1})$, where $W_{0}$ and $h_{0}$ are initial half width and half height of the slab, respectively. The factor $\alpha$ and another factor of linearity must then be found empirically from experiments, with a proviso that each resulting equation only gives reasonable lateral spread prediction within the ranges of conditions for which they were empirically determined (chitkara1966some). For example, for low and high carbon steel kazeminezhad2005width wrote equation (1) as
@@
Among a wide range of empirical formulations suggested for predicting lateral spread (chitkara1966some), the one proposed by kobayashi1978influence is often referred to as being reliable for round wire (utsunomiya2001three; kazeminezhad2008error; vallellano2008analysis). utsunomiya2001three and vallellano2008analysis attribute the following equation to kobayashi1978influence, and we will refer to it here by this name:
@@
Finite element (FE) analysis has provided valuable insights into wire rolling, similar to other metal-forming processes. vallellano2008analysis performed a 3D numerical analysis using the Abaqus finite element software to study contact stress distributions, residual stresses, and lateral spread. While 3D FE analysis offers highly detailed information, it is computationally expensive, especially when it is used in design iterations where several simulations are required. Although some studies have investigated FE simulations of wire rolling (carlsson1998contact; vallellano2008analysis; hwang2021influence), computational cost data are generally not reported. However, to give an indication, a 2D FE simulation of half-of-the-thickness sheet rolling required 15.9 CPU hours for a reliable case (30 elements through the half-thickness) and 0.25 CPU hours for a coarse case (5 elements) (flanagan2025); since 3D wire-rolling simulations are considerably more complex, they are expected to be significantly more expensive yet. In addition to high computational cost, FE results require validation against a reliable reference. Whether the objective is process design or FE validation, a review of the literature further reveals a notable gap in the availability of models with well-defined, traceable assumptions for the rolling of round wire.
@@
Using similar techniques to the asymptotic mathematical modelling of sheet rolling (erfanian2025), here an asymptotic mathematical model is developed for the regime of a small friction coefficient and large length-to-thickness ratio, consistent with the slab method of kazeminezhad2006pressure. The further assumption that the width is comparably smaller than the length of wire in the roll gap simplifies the problem to plane-stress, where the stress in $z$ direction is negligible. This allows for finding the lateral spread without the need for fitting parameters or solving a problem in a complicated 3D stress state. Even so, the derivation of equations in all directions and the use of asymptotic analysis provide a foundation for further improvements to the model in the future, unlike the slab method, which remains inherently limited.
@@
An outline of the paper is as follows. The mathematical model is explained in Section 2, beginning with section 2.1 introducing the simplifying assumptions. The scaling parameters and resulting non-dimensionalised governing equations are given in Section 2.2 for a rigid perfectly plastic material, and their solution is derived in Section 2.3. Section 2.4 summarises the solutions and outlines the computational methods used to evaluate them. The experimental procedure for obtaining data is detailed in Section 3, which is used to validate the lateral spread predictions of the model in Section 4.1, alongside FE data from vallellano2008analysis. The Longitudinal spread is compared against experimental data in Section 4.2 and the roll pressure predictions are validated against FE data from vallellano2008analysis and presented in Section 4.3. Having validated the model, a parametric study is presented in Section 5, exploring both the range of validity of the model in Section 5.1 and the influence of parameters including friction, reduction, and roll diameter on the lateral spread in Section 5.2. Finally, a discussion and potential directions for future research are provided in Section 6.
@@ 2.2 Scalings and the non-dimensionalised governing equations
Asymptotic (or perturbation) methods (hinch) are widely used in applied mechanics, exploiting the smallness of an inherent parameter to obtain approximate analytical/semi-analytical solutions with controllable errors. The first step in applying this technique is to rescale the governing equations and parameters based on the inherent scales of the problem. This process removes physical units and expresses all quantities in a comparable dimensionless form, allowing their relative magnitudes to be evaluated. The resulting solution can then be redimensionalized using the same scales to provide an engineering solution in physical units. Here, the horizontal distance is scaled with the characteristic roll-gap length, $\hat{\ell}$, so that $x=\hat{x}$$\hat{\ell}$ gives a dimensionless horizontal coordinate $x$, and the dimensional horizontal position $\hat{x}$ can be recovered from $x$ using $\hat{x}=\hat{\ell}x$. Similarly, the vertical distance is scaled with the initial sheet half-thickness, $\hat{h}_{\!\>\!\mathrm{A}}$, so that $y=\hat{y}/\hat{h}_{\!\>\!\mathrm{A}}$. For circular rolls of radius $\hat{R}$, the roll-gap length $\hat{\ell}$ is equal to
@@
$=\mp\delta\beta\!\left(\!{\sigma}_{yy}-2\delta^{2}\frac{\mathrm{d}h}{\mathrm{d}x}{\sigma}_{xy}+\delta^{2}\!\left(\frac{\mathrm{d}h}{\mathrm{d}x}\right)^{\!\!2}\!{\sigma}_{xx}\!\right)\hfil\text{on }y=h(x).$
@@
where $h_{\text{out}}$ and $b_{\text{out}}$ are respectively half of the final thickness and width, and $h_{A}=1$ and $b_{A}$ are respectively half of the initial thickness and width at location A. Vertical symmetry requires the boundary conditions
@@
$\int_{-b(x)}^{b(x)}\int_{-h(x)}^{h(x)}u\ \mathrm{d}y\mathrm{d}z=4b_{A},$ (21)
where $h=u=1$ and $b=b_{A}$ at point A gives the constant of integration as $4b_{A}$. Equation (21) effectively says that, since no material is lost or gained during rolling and the material is incompressible, the volume rate of flow across the wire cross-section is a constant independent of where along the wire it is measured at.
@@ 2.3 Leading-order asymptotic solution
Following non-dimensionalisation, all terms in the governing equations are now the same size except for the explicit factors of $\delta$, which is small. Following the technique of asymptotics (see, e.g. hinch), we now expand in an asymptotic series in powers of the small parameter $\delta$, such that each successive term is asymptotically smaller than the preceding one. Whilst many complicated expansions are sometimes needed, here a simple expansion in powers of $\delta$ is sufficient, as is the case for sheet rolling (erfanian2025); this is suggested from the fact that the small parameter $\delta$ only appears as integer powers of $\delta$ in the governing equations and boundary conditions above, and may be verified a posteriori once the solution has been calculated. We therefore expand any unknown variable as
@@
with $\phi^{(0)}$ not identically zero so that $\phi=O(1)$. Here, $\phi$ represents any of the unknown variables $u$, $v$, $w$, $\sigma_{ij}$, $s_{ij}$, $p$ and $\lambda$; for example, $\sigma_{xx}=\sigma^{(0)}_{xx}+\delta\sigma^{(1)}_{xx}+\delta^{2}\sigma^{(2)}_{xx}$$\cdots$. Substituting the asymptotic expansion into the governing equations (13) to (21) allows smaller terms to be neglected relative to dominant ones successively, establishing an ordered hierarchy that leads to an approximate but systematically improvable solution with a quantifiable error. While this formalism could be used to compute the solution to an arbitrary order of accuracy (for example, erfanian2025 consider up to the second order terms $\sigma^{(2)}_{xx}$ for sheet rolling), here we will only consider the leading order terms. We therefore neglect all terms of $O(\delta)$ or higher to derive the leading-order equations. It should be noted that neglecting $O(\delta)$ terms in the expansion does not mean we are setting $\delta$ to zero, but rather that we are calculating the leading-order solution which represents the dominant physical behaviour for small, yet realistic, values of $\delta$, with a relative error of order $O(\delta)$ which could subsequently be corrected for by including higher-order terms. From the hydrostatic pressure (15) at leading order
@@
This is the same equation as was derived using the slab method by kazeminezhad2006pressure, although our formulae for $\sigma^{(0)}_{xx}$ and $\sigma^{(0)}_{yy}$ here are different. Equation (27) provides an ordinary differential equation (ODE) in terms of $p^{(0)}$ and $b(x)$, when combined with stress solutions (24).
@@
$\lambda^{(0)}=\frac{b_{A}\mathrm{d}h/\mathrm{d}x}{bh^{2}s^{(0)}_{yy}}$, (30)
@@
$\lambda^{(0)}=\frac{b_{A}\mathrm{d}b/\mathrm{d}x}{b^{2}hp^{(0)}}.$ (31)
@@
$4b_{A}\sigma^{(0)}_{xx}(x=\mathrm{A})$ $=F_{\mathrm{A}}$ at point A, (35a)
@@ 2.4 Summary and numerical evaluation
The solution for $p^{(0)}$ is chosen to satisfy the forward and backward tension conditions, which are taken to be zero for the results presented below. Therefore, $p^{(0)}$ is solved by integrating equation (34) forward from the entrance with positive sign of friction, and integrating equation (34) backwards from the exit with positive sign of friction, using the Matlab ODE solver ode45 (MATLAB:R2024a_u1). This is the same solution as the slab method, and the two curves thus produced are referred to as the pressure hill, and the intersection determines the location of the neutral point.
After solving for stresses, $b(x)$ and consequently the lateral spread, $W_{t}$, is determined by integrating either equation (32) or (33) from point A to the exit. The integration is performed using the Matlab ODE solver ode45, with the initial condition $b(0)=b_{A}$. For the rolling of a cylindrical wire, we assume $b_{A}=h_{A}=1$, as stated in section 2.1.1 and further justified in A, and consistent with the non-dimensionalisation. Nevertheless, retaining $b_{A}$ explicitly in the governing equations allows the same model to be applied to subsequent roll stands, where the initial cross-section is no longer circular but approximately rectangular with bulged edges. In such cases, location $A$ corresponds to the roll-gap entrance. As detailed in the introduction and shown in figure 2, the area of the rectangular cross-section is assumed to be equal to that of the actual wire in the roll gap. As a result, the lateral spread is determined by equating the cross-sectional area of the real shape to that of the rectangle, with $\hat{W}_{c}$ the width in contact with the roll and $\hat{W}_{t}$ the total lateral spread then given by (5) and (6) respectively.

### v3

@@ 1 Introduction
Applications including sawblades, springs, piston rings, and transformers depend critically on flattened wires (Utsunomiya et al., 2001). These wires are usually made by a flat rolling technique, where a wire with a circular cross-section is cold rolled between cylindrical rolls — sometimes through several passes — to achieve a particular width and thickness (as illustrated in figure 1).
Figure 1: A diagram of wire flat rolling process; the wire initially has a circular cross-section with diameter $d_{0}$ and is flattened to a barrel shape cross section with the lateral spread $W_{t}$ and the contact width $W_{c}$. Adapted from Figure 1 of Carlsson (1998).
@@
In cold rolling of a sheet, deformation in the lateral direction often remains within the elastic range due to the product’s geometry. However, this is not the case when an initial width-to-thickness ratio is less than 6, such as in wide strip or plate, or as low as 1 in round wire (Chitkara and Johnson, 1966). In such cases, the plastic flow in the roll gap at the start of rolling schedule is inherently three-dimensional, complicating the analysis. For round wire, this complexity is further increased by the transformation of a round cross-section into a rectangular shape with bulged edges during the first pass.
Although research exists on the modelling of lateral spread in thick plate rolling (e.g. Chitkara and Johnson, 1966; Lahoti and Kobayashi, 1974; Oh and Kobayashi, 1975; El-Nikhaily, 1979; Kennedy, 1987), wire flat rolling has received less attention. Kazeminezhad et al. (2005; 2006; 2008) published a series of papers studying different parameters in wire rolling. Kazeminezhad and Karimi Taheri (2005) developed a relationship for the width of the contact area between the rolls and the wire. Their equation was based on experimental and FE observations of the formation of X-shaped shear bands in the wire cross-section during the rolling process (Semiatin and Jonas, 1984; Pesin et al., 2002), along with the assumption that as the height reduction increases, the shear bands rotate while maintaining a constant length. Kazeminezhad and Karimi Taheri (2006) then applied this equation alongside the slab method, leading to a pressure hill distribution. The study employs the Tresca yield criterion, assuming the longitudinal stress $\sigma_{xx}$ to be the minimum stress, the vertical stress $\sigma_{yy}$ to be the maximum stress, and the compressive stress in the lateral direction, $\sigma_{zz}$, to lie between them. Although $\sigma_{zz}$ is implicitly included in the yield function, and therefore in calculating the roll pressure, the value of $\sigma_{zz}$ is unknown and must be approximated to find the lateral spread. Kazeminezhad and Karimi Taheri (2005) therefore assumed that $\sigma_{zz}$ takes a value between plane strain and plane stress, written as $\alpha\sigma_{yy}$, where $\alpha$ is a fitting parameter taking values between 0 and 0.5. Using flow rule equations and setting $\sigma_{xx}$ to be zero at the end of the roll gap, they then wrote
@@
where $W_{t}$ is the lateral spread, $d_{0}$ is the initial diameter, and $2h_{1}$ is the final height of the wire. The first use of this formulation has been attributed to Hill (1955) for plate rolling and has also been derived by several other authors (e.g. Chitkara and Johnson, 1966; Applied Mechanics Group and Sparling, 1961), who experimentally observed a linear relationship between $\ln(W_{t}/W_{0})$ and $\ln(h_{0}/h_{1})$, where $W_{0}$ and $h_{0}$ are initial half width and half height of the slab, respectively. The factor $\alpha$ and another factor of linearity must then be found empirically from experiments, with a proviso that each resulting equation only gives reasonable lateral spread prediction within the ranges of conditions for which they were empirically determined (Chitkara and Johnson, 1966). For example, for low and high carbon steel Kazeminezhad and Karimi Taheri (2005) wrote equation (1) as
@@
Among a wide range of empirical formulations suggested for predicting lateral spread (Chitkara and Johnson, 1966), the one proposed by Kobayashi (1978) is often referred to as being reliable for round wire (Utsunomiya et al., 2001; Kazeminezhad et al., 2008; Vallellano et al., 2008). Utsunomiya et al. (2001) and Vallellano et al. (2008) attribute the following equation to Kobayashi (1978), and we will refer to it here by this name:
@@
Finite element (FE) analysis has provided valuable insights into wire rolling, similar to other metal-forming processes. Vallellano et al. (2008) performed a 3D numerical analysis using the Abaqus finite element software to study contact stress distributions, residual stresses, and lateral spread. While 3D FE analysis offers highly detailed information, it is computationally expensive, especially when it is used in design iterations where several simulations are required. Although some studies have investigated FE simulations of wire rolling (Carlsson, 1998; Vallellano et al., 2008; Hwang et al., 2021), computational cost data are generally not reported. However, to give an indication, a 2D FE simulation of half-of-the-thickness sheet rolling required 15.9 CPU hours for a reliable case (30 elements through the half-thickness) and 0.25 CPU hours for a coarse case (5 elements) (Flanagan et al., 2025); since 3D wire-rolling simulations are considerably more complex, they are expected to be significantly more expensive yet. In addition to high computational cost, FE results require validation against a reliable reference. Whether the objective is process design or FE validation, a review of the literature further reveals a notable gap in the availability of models with well-defined, traceable assumptions for the rolling of round wire.
@@
Using similar techniques to the asymptotic mathematical modelling of sheet rolling (Erfanian et al., 2025), here an asymptotic mathematical model is developed for the regime of a small friction coefficient and large length-to-thickness ratio, consistent with the slab method of Kazeminezhad and Karimi Taheri (2006). The further assumption that the width is comparably smaller than the length of wire in the roll gap simplifies the problem to plane-stress, where the stress in $z$ direction is negligible. This allows for finding the lateral spread without the need for fitting parameters or solving a problem in a complicated 3D stress state. Even so, the derivation of equations in all directions and the use of asymptotic analysis provide a foundation for further improvements to the model in the future, unlike the slab method, which remains inherently limited.
@@
An outline of the paper is as follows. The mathematical model is explained in Section 2, beginning with section 2.1 introducing the simplifying assumptions. The scaling parameters and resulting non-dimensionalised governing equations are given in Section 2.2 for a rigid perfectly plastic material, and their solution is derived in Section 2.3. Section 2.4 summarises the solutions and outlines the computational methods used to evaluate them. The experimental procedure for obtaining data is detailed in Section 3, which is used to validate the lateral spread predictions of the model in Section 4.1, alongside FE data from Vallellano et al. (2008). The Longitudinal spread is compared against experimental data in Section 4.2 and the roll pressure predictions are validated against FE data from Vallellano et al. (2008) and presented in Section 4.3. Having validated the model, a parametric study is presented in Section 5, exploring both the range of validity of the model in Section 5.1 and the influence of parameters including friction, reduction, and roll diameter on the lateral spread in Section 5.2. Finally, a discussion and potential directions for future research are provided in Section 6.
@@ 2.2 Scalings and the non-dimensionalised governing equations
Asymptotic (or perturbation) methods (Hinch, 1991) are widely used in applied mechanics, exploiting the smallness of an inherent parameter to obtain approximate analytical/semi-analytical solutions with controllable errors. The first step in applying this technique is to rescale the governing equations and parameters based on the inherent scales of the problem. This process removes physical units and expresses all quantities in a comparable dimensionless form, allowing their relative magnitudes to be evaluated. The resulting solution can then be redimensionalized using the same scales to provide an engineering solution in physical units. Here, the horizontal distance is scaled with the characteristic roll-gap length, $\hat{\ell}$, so that $x=\hat{x}$$\hat{\ell}$ gives a dimensionless horizontal coordinate $x$, and the dimensional horizontal position $\hat{x}$ can be recovered from $x$ using $\hat{x}=\hat{\ell}x$. Similarly, the vertical distance is scaled with the initial sheet half-thickness, $\hat{h}_{\!\>\!\mathrm{A}}$, so that $y=\hat{y}/\hat{h}_{\!\>\!\mathrm{A}}$. For circular rolls of radius $\hat{R}$, the roll-gap length $\hat{\ell}$ is equal to
@@
$=\mp\delta\beta\!\left(\!{\sigma}_{yy}-2\delta^{2}\frac{\mathrm{d}h}{\mathrm{d}x}{\sigma}_{xy}+\delta^{2}\!\left(\frac{\mathrm{d}h}{\mathrm{d}x}\right)^{\!\!2}\!{\sigma}_{xx}\!\right)\quad\text{on }y=h(x).$
@@
where $h_{\text{out}}$ and $b_{\text{out}}$ are respectively half of the final thickness and width, and $h_{\!\>\!\mathrm{A}}=1$ and $b_{\!\>\!\mathrm{A}}$ are respectively half of the initial thickness and width at location A. Vertical symmetry requires the boundary conditions
@@
$\int_{-b(x)}^{b(x)}\int_{-h(x)}^{h(x)}u\ \mathrm{d}y\mathrm{d}z=4b_{\!\>\!\mathrm{A}},$ (21)
where $h=u=1$ and $b=b_{\!\>\!\mathrm{A}}$ at point A gives the constant of integration as $4b_{\!\>\!\mathrm{A}}$. Equation (21) effectively says that, since no material is lost or gained during rolling and the material is incompressible, the volume rate of flow across the wire cross-section is a constant independent of where along the wire it is measured at.
@@ 2.3 Leading-order asymptotic solution
Following non-dimensionalisation, all terms in the governing equations are now the same size except for the explicit factors of $\delta$, which is small. Following the technique of asymptotics (see, e.g. Hinch, 1991), we now expand in an asymptotic series in powers of the small parameter $\delta$, such that each successive term is asymptotically smaller than the preceding one. Whilst many complicated expansions are sometimes needed, here a simple expansion in powers of $\delta$ is sufficient, as is the case for sheet rolling (Erfanian et al., 2025); this is suggested from the fact that the small parameter $\delta$ only appears as integer powers of $\delta$ in the governing equations and boundary conditions above, and may be verified a posteriori once the solution has been calculated. We therefore expand any unknown variable as
@@
with $\phi^{(0)}$ not identically zero so that $\phi=O(1)$. Here, $\phi$ represents any of the unknown variables $u$, $v$, $w$, $\sigma_{ij}$, $s_{ij}$, $p$ and $\lambda$; for example, $\sigma_{xx}=\sigma^{(0)}_{xx}+\delta\sigma^{(1)}_{xx}+\delta^{2}\sigma^{(2)}_{xx}$$\cdots$. Substituting the asymptotic expansion into the governing equations (13) to (21) allows smaller terms to be neglected relative to dominant ones successively, establishing an ordered hierarchy that leads to an approximate but systematically improvable solution with a quantifiable error. While this formalism could be used to compute the solution to an arbitrary order of accuracy (for example, Erfanian et al. (2025) consider up to the second order terms $\sigma^{(2)}_{xx}$ for sheet rolling), here we will only consider the leading order terms. We therefore neglect all terms of $O(\delta)$ or higher to derive the leading-order equations. It should be noted that neglecting $O(\delta)$ terms in the expansion does not mean we are setting $\delta$ to zero, but rather that we are calculating the leading-order solution which represents the dominant physical behaviour for small, yet realistic, values of $\delta$, with a relative error of order $O(\delta)$ which could subsequently be corrected for by including higher-order terms. From the hydrostatic pressure (15) at leading order
@@
This is the same equation as was derived using the slab method by Kazeminezhad and Karimi Taheri (2006), although our formulae for $\sigma^{(0)}_{xx}$ and $\sigma^{(0)}_{yy}$ here are different. Equation (27) provides an ordinary differential equation (ODE) in terms of $p^{(0)}$ and $b(x)$, when combined with stress solutions (24).
@@
$\lambda^{(0)}=\frac{b_{\!\>\!\mathrm{A}}\mathrm{d}h/\mathrm{d}x}{bh^{2}s^{(0)}_{yy}}$, (30)
@@
$\lambda^{(0)}=\frac{b_{\!\>\!\mathrm{A}}\mathrm{d}b/\mathrm{d}x}{b^{2}hp^{(0)}}.$ (31)
@@
$4b_{\!\>\!\mathrm{A}}\sigma^{(0)}_{xx}(x=\mathrm{A})$ $=F_{\mathrm{A}}$ at point A, (35a)
@@ 2.4 Summary and numerical evaluation
The solution for $p^{(0)}$ is chosen to satisfy the forward and backward tension conditions, which are taken to be zero for the results presented below. Therefore, $p^{(0)}$ is solved by integrating equation (34) forward from the entrance with positive sign of friction, and integrating equation (34) backwards from the exit with positive sign of friction, using the Matlab ODE solver ode45 (MathWorks, 2024). This is the same solution as the slab method, and the two curves thus produced are referred to as the pressure hill, and the intersection determines the location of the neutral point.
After solving for stresses, $b(x)$ and consequently the lateral spread, $W_{t}$, is determined by integrating either equation (32) or (33) from point A to the exit. The integration is performed using the Matlab ODE solver ode45, with the initial condition $b(0)=b_{\!\>\!\mathrm{A}}$. For the rolling of a cylindrical wire, we assume $b_{\!\>\!\mathrm{A}}=h_{\!\>\!\mathrm{A}}=1$, as stated in section 2.1.1 and further justified in A, and consistent with the non-dimensionalisation. Nevertheless, retaining $b_{\!\>\!\mathrm{A}}$ explicitly in the governing equations allows the same model to be applied to subsequent roll stands, where the initial cross-section is no longer circular but approximately rectangular with bulged edges. In such cases, location $A$ corresponds to the roll-gap entrance. As detailed in the introduction and shown in figure 2, the area of the rectangular cross-section is assumed to be equal to that of the actual wire in the roll gap. As a result, the lateral spread is determined by equating the cross-sectional area of the real shape to that of the rectangle, with $\hat{W}_{c}$ the width in contact with the roll and $\hat{W}_{t}$ the total lateral spread then given by (5) and (6) respectively.